### v1

# Semantics-Oriented Multitask Learning for DeepFake Detection: A Joint Embedding Approach

Abstract
In recent years, the multimedia forensics and security community has seen remarkable progress in multitask learning for DeepFake (i.e., face forgery) detection. The prevailing strategy has been to frame DeepFake detection as a binary classification problem augmented by manipulation-oriented auxiliary tasks. This strategy focuses on learning features specific to face manipulations, which exhibit limited generalizability. In this paper, we delve deeper into semantics-oriented multitask learning for DeepFake detection, leveraging the relationships among face semantics via joint embedding. We first propose an automatic dataset expansion technique that broadens current face forgery datasets to support semantics-oriented DeepFake detection tasks at both the global face attribute and local face region levels. Furthermore, we resort to joint embedding of face images and their corresponding labels (depicted by textual descriptions) for prediction. This approach eliminates the need for manually setting task-agnostic and task-specific parameters typically required when predicting labels directly from images. In addition, we employ a bi-level optimization strategy to dynamically balance the fidelity loss weightings of various tasks, making the training process fully automated. Extensive experiments on six DeepFake datasets show that our method improves the generalizability of DeepFake detection and, meanwhile, renders some degree of model interpretation by providing human-understandable explanations.
Index Terms:
DeepFake detection, face semantics, multitask learning, joint embedding.
I Introduction
The emergence of deep learning models [1, 2, 3, 4] has greatly simplified and automated the production of realistic counterfeit face images and videos, commonly referred to as DeepFakes. This seriously threatens the authenticity and integrity of digital visual media. In response, a wide array of DeepFake detection methods have been developed to efficiently screen manipulated face imagery.
The early DeepFake detectors [5, 6, 7] were largely shaped by established photo forensics techniques [8], through the analysis of statistical anomalies [9], visual artifacts [10, 11, 12], and physical inconsistencies [13, 14, 15]. The advent of deep learning has spurred the development of numerous learning-based detection methods [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37], which operate as binary classifiers to predict whether a test image is real or fake. This formulation can be augmented by incorporating auxiliary tasks, such as manipulation type identification [38, 39], manipulation parameter estimation [40], blending boundary detection [21], and face reconstruction [23, 26]), leading to a multitask learning paradigm [20, 21, 22, 23, 24, 25, 26, 27]. The underlying assumption is that the primary task will likely gain from including auxiliary tasks through knowledge transfer. However, most auxiliary tasks are specific to manipulations, which may hinder the ability to generalize. Recently, Zou et al. [41] introduced a semantics-oriented multitask learning paradigm for DeepFake detection. They defined face forgery as the use of computational methods that modify semantic face attributes beyond the thresholds of human discrimination. Although conceptually and computationally appealing, training semantics-oriented DeepFake detectors would require considerable human annotations to specify manipulation degree parameters and semantic label hierarchies [41]. Moreover, the allocation of task-agnostic and task-specific parameters is optimized manually in multitask learning, which is bound to be suboptimal in facilitating knowledge transfer among tasks.
Figure 1: Illustration of the label hierarchy of the expanded FF++ dataset [18] by the proposed semantics-oriented dataset expansion technique.
In this paper, we further pursue the promising direction of semantics-oriented multitask learning for DeepFake detection with two important improvements. First, we present a dataset expansion technique that automatically enriches the binary labeling of existing DeepFake datasets to semantic hierarchical labeling while increasing their sizes (see Fig. 1). A crucial element of our technique is to incorporate the physical_inconsistency node that encompasses local face region manipulations, without manually identifying interdependent changes in global face attributes (if any). Second, unlike discriminative learning that directly predicts target labels from images, we take inspiration from the recent advances in vision-language correspondence [42] and encode both the input image and the target label hierarchy (represented by text templates) into a shared feature space. This enables us to view model parameter sharing/splitting in multitask learning from the new perspective of model capacity allocation. That is, we share the image encoder parameters across all tasks, with the capacity allocated to individual tasks being end-to-end optimized. In addition, we adopt bi-level optimization [43] to minimize the primary task with respect to the fidelity loss weightings [44] in the validation set as a way of task prioritization. The parameters of the DeepFake detector are learned by minimizing the weighted summation of fidelity losses for individual tasks.
We demonstrate the effectiveness of our Semantics-oriented Joint Embedding DeepFake Detector (SJEDD) through extensive experiments on six datasets [18, 45, 46, 47, 48, 41]. SJEDD consistently improves the cross-dataset, cross-manipulation, and cross-attribute detection performance compared to eighteen state-of-the-art DeepFake detectors.
II Related Work
In this section, we provide a concise review of DeepFake detectors, emphasizing the multitask learning approach. We also discuss joint embedding architectures in the broader context of machine learning.
II-A DeepFake Detectors
Non-learning-based detectors rely on extensive human expertise for their design. The first type of knowledge involves mathematical models of the image source, which summarizes the essential elements of what constitutes an authentic image [9]. The second type of knowledge involves computational models of the digital photography pipeline, along which each step can be exploited to expose forgeries [49], including the lens distortion and flare [10, 11], color filter array [50], sensor noise [12], pixel response non-uniformity [51], camera response function [52], and file storage format [53]. The third type of knowledge involves physical and geometric models of the three-dimensional world, especially those that capture the interactions of light, optics, and objects. The image forgeries can be exposed by detecting physical and geometric inconsistencies, such as those found in illumination [13], shadow [14], reflection [15], and motion [54]. Generally, non-learning-based methods are limited by the precision and expressiveness of manually crafted models and thus achieve limited success.
Learning-based methods detect DeepFake images and videos by learning directly from data. Afchar et al. [17] proposed to examine DeepFake videos at the mesoscopic level. Zhou et al. [16] designed a two-stream network to jointly analyze camera characteristics and assess visual artifacts. Li and Lyu [55] aimed to detect affine warping artifacts, while Li et al. [21] focused on locating image blending boundaries in the intermediate steps of face manipulations. Dong et al. [25] revealed a shortcut termed “implicit identity leakage,” and described an identity-unaware DeepFake detector. Gu et al. [56] and Zheng et al. [57] examined spatiotemporal inconsistencies at the signal level, while Haliassos et al. [28] did so at the semantic level through lipreading pretraining. Xu et al. [58] proposed a thumbnail layout to trade spatial modeling for temporal modeling. Another line of research that has received significant attention involves detecting abnormal physiological reactions from blinking of the eyes [59], pupil shape [60], head pose [61], corneal specularity [62], and action units [63]. Complex architectures, such as capsule networks [64] and vision Transformers [65, 31], along with advanced training strategies such as adversarial training [24], test-time training [66], graph learning [34, 36], contrastive learning [32, 37], and out-of-distribution detection [67], have also been used for DeepFake detection.
Multitask learning-based methods have emerged as a promising offshoot of learning-based methods for DeepFake detection in the past five years. Notable auxiliary tasks include manipulation localization [20, 24] (with blending boundary detection [21] as a special case), manipulation classification and attribution [38, 39], manipulation parameter estimation [40], face image reconstruction [23, 26], and tasks that promote robust and discriminative feature learning [27]. Most auxiliary tasks are manipulation-oriented, which may increase the risk of overfitting and result in poor generalization to novel manipulations. This work improves upon the recently proposed semantics-oriented multitask learning paradigm [41]. From the data perspective, we introduce a dataset expansion technique that not only increases the sizes of existing DeepFake datasets but also enriches them with a collection of labels organized in a semantic hierarchical graph (see Fig. 1). From the model perspective, we leverage joint embedding instead of discriminative learning to automate model parameter sharing/splitting and to facilitate task relationship modeling. Concurrently, Sun et al. [33] proposed VLFFD, which also implemented DeepFake detection via joint embedding. The proposed SJEDD differs significantly from VLFFD in terms of motivation (i.e., enhancing semantics-oriented detection as opposed to leveraging multimodal data), task nature (i.e., semantics-oriented tasks rather than manipulation-oriented ones), and model optimization (i.e., principled bi-level optimization instead of heuristic coarse-to-fine training). Consequently, under the same experimental setups, the proposed SJEDD outperforms VLFFD on various DeepFake datasets by wide margins (see Sec. V-B).
Figure 2: Pipeline for local face region manipulation. When searching for the best candidate as the source image, we minimize the Euclidean distance between $68$ detected landmarks [21] of the target and candidate face images (excluding those with the same identity).
II-B Joint Embedding Architectures
Joint embedding architectures [68] aim to learn embeddings that are close for compatible inputs (e.g., an input image and its target label) and distant for incompatible inputs, as opposed to discriminative approaches that directly predict labels from inputs. The idea can be traced back to the Siamese network [69], initially proposed for signature verification and later adapted to computer vision applications like face verification [70] and recognition [71], and video feature learning [72]. With the rise of self-supervised learning, a thorough exploration of joint embedding architectures is conducted, contrastively (e.g., MoCo [73] and SimCLR [74]) and non-contrastively (e.g., Barlow Twins [75] and I-JEPA [76]). The more recent vision-language models [77, 42] are also manifestations of joint embedding architectures and are quickly adapted to multiple downstream tasks such as image quality assessment [78], realistic image enhancement [79], image-text retrieval [80], and multimodal fake news detection [81]. Here, we use CLIP [42], a vision-language model pre-trained on massive image-text pairs, to implement SJEDD.
III Semantics-Oriented Dataset Expansion: A Demonstration on FF++
In this section, we describe our dataset expansion method in detail, consisting of two steps: data augmentation and data relabeling. Data augmentation generates new images by manipulating local face regions. Data relabeling automatically assigns DeepFake images with a set of labels through semantic contextualization of face manipulations [41]. To demonstrate the proposed dataset expansion technique, we apply it to expand the widely used FF++ dataset [18], while acknowledging that our method can also be applied to other DeepFake datasets.
III-A Data Augmentation
We augment FF++ by manipulating four local face regions, i.e., eye, lip, mouth, and nose, as illustrated in Fig. 2. Given a real face as the target, we initially search for the best source face (real or fake) by minimizing the Euclidean distance between their corresponding $68$ detected landmarks [21] while excluding faces with the same identity [21]. Next, we linearly blend the source local region, indicated by a Gaussian-blurred region-of-interest mask, into the target image, followed by color correction [82].
III-B Data Relabeling
Inspired by [41], we first construct a semantic hierarchical graph to contextualize the four face manipulations in FF++: Deepfakes [1], FaceSwap [83], Face2Face [84], and NeuralTextures [4]. The former two manipulations swap faces between images to achieve identity manipulation, while the latter two manipulations change the expression of the target face through manipulation of the mouth region. Thus, we instantiate two global face attribute nodes: expression and identity. We also include six leaf nodes, representing nonoverlapping local face regions: eye, eyebrow, lip, mouth, nose, and skin. For the two identity manipulations based on face swapping, all local face regions are assumed to be altered. In contrast, for the two expression manipulations, only lip and mouth regions are affected. To accommodate local face manipulations in the graph, we introduce an additional global face attribute node—physical_inconsistency—which signifies images where a local face region is replaced by the corresponding region from another image (real or fake), resulting in physical incompatibility between this local region and the other real face regions.
TABLE I: Summary of the expanded FF++ dataset by the proposed dataset expansion technique, where “1” means the corresponding face attribute (or region) is modified and “0” otherwise. “N/A” indicates the manipulation status is not accessible (i.e., unobserved)
Manipulation / Region(s) Face Expression Identity Physical_inconsistency Eye Eyebrow Lip Mouth Nose Skin
Data Augmentation Eye 1 N/A 0 1 1 0 0 0 0 0
Lip and Mouth 1 N/A 0 1 0 0 1 1 0 0
Nose 1 0 0 1 0 0 0 0 1 0
Data Relableling Deepfakes [1] 1 N/A 1 1 1 1 1 1 1 1
Face2Face [84] 1 1 0 1 0 0 1 1 0 0
FaceSwap [83] 1 N/A 1 1 1 1 1 1 1 1
NeuralTextures [4] 1 1 0 1 0 0 1 1 0 0
III-C Discussion
The proposed dataset expansion technique allows some labels in the semantic hierarchical graph to remain unobserved, which helps encourage labeling reliability. For instance, in the case of face swapping, it is unclear whether the expression has been altered, so we leave this label unobserved. Meanwhile, in the expanded dataset, a single face manipulation can simultaneously alter multiple face attributes (e.g., identity and physical_consistency by FaceSwap [83]), while different manipulations can also modify the same attribute (e.g., identity by Deepfakes [1] and FaceSwap). That is, our technique emphasizes a semantics-oriented rather than manipulation-oriented contextualization of face manipulations, akin to FFSC [41], but achieved without expensive human labeling. Table I presents the list of face manipulations in the expanded FF++ dataset, along with the corresponding modified face attributes and regions.
IV Proposed Method: SJEDD
In this section, we describe Semantics-oriented Joint Embedding DeepFake Detector (SJEDD) in detail, our joint embedding approach for semantics-oriented multitask DeepFake detection, including 1) general problem formulation, 2) joint embedding for detection, and 3) bi-level optimization for parameter estimation. The overall structure of SJEDD is shown in Fig. 3.
IV-A Problem Formulation
Given an input face image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, we assign to it a vector of binary labels $\bm{y}=\{0,1\}^{C}$, indicating one primary and $C-1$ auxiliary tasks. Specifically, $y_{1}=1$ signifies a forged face, and $y_{1}=0$ means the opposite. For $i>1$, $y_{i}=1$ indicates that the $i$-th face attribute/region is fake, and $y_{i}=0$ otherwise. Next, we define an unnormalized probability distribution over the label hierarchy [85]:
$\begin{split}\tilde{p}(\bm{y}|\bm{x})=&\prod_{i}e^{s_{i}(\bm{x})\mathbb{I}[y_{i}=1]}\prod_{j,i\in\mathcal{P}_{j}}\mathbb{I}\left[\left(\sum_{i}y_{i},y_{j}\right)\neq(0,1)\right]\\ &\prod_{i,j\in\mathcal{C}_{i}}\mathbb{I}\left[\left(y_{i},\sum_{j}y_{j}\right)\neq(1,0)\right],\end{split}$ (1)
where $s_{i}(\bm{x})$ represents the raw score of the detector for the $i$-th label, and $\mathbb{I}[\cdot]$ is an indicator function. Eq. (1) encodes two label relations: 1) if the $j$-th node is activated (i.e., $y_{j}=1$), at least one of its parent nodes, collectively indexed by $\mathcal{P}_{j}$, must also be activated; and 2) if the $i$-th node is activated, at least one of its child nodes, indexed by $\mathcal{C}_{i}$, must also be activated. We then normalize Eq. (1) to obtain the joint probability
$\hat{p}(\bm{y}|\bm{x})=\frac{\tilde{p}(\bm{y}|\bm{x})}{Z(\bm{x})},\,\textrm{where}\,Z(\bm{x})=\sum_{\bm{y}^{\prime}}\tilde{p}(\bm{y}^{\prime}|\bm{x}).$ (2)
Due to the small scale of our label hierarchy, $Z(\bm{x})$ can be easily computed using matrix multiplication. We further compute the marginal probability of a given label $y_{i}$ by marginalizing the joint probability over all other labels:
$\hat{p}(y_{i}|\bm{x})=\sum_{\bm{y}\setminus y_{i}}\hat{p}(\bm{y}|\bm{x}),$ (3)
where $\setminus$ denotes the set difference operation.
IV-B Joint Embedding
We first introduce text templates to encode the label hierarchy and then implement joint embedding using vision-language correspondence to estimate the raw scores for each label.
Figure 3: System diagram of SJEDD.
IV-B1 Text Templates
For the root node face, the text template is defined as “a photo of a fake face.” At the global face attribute level, we use the template “A photo of a face with the global attribute of $\{g\}$ altered,” where $g$ belongs to the set of global face attributes {expression, identity, physical_inconsistency}. At the level of the local face region, the template is defined as “A photo of a face with the local $\{l\}$ region altered,” where $l$ is in the set of local face regions {eye, eyebrow, lip, mouth, nose, skin}.
IV-B2 Vision-Language Correspondence
We construct two encoders: an image encoder $\bm{f}_{\bm{\phi}}:\mathbb{R}^{H\times W\times 3}\mapsto\mathbb{R}^{N\times 1}$, parameterized by $\bm{\phi}$, which computes the visual embedding $\bm{f}_{\bm{\phi}}(\bm{x})$ for an input face image $\bm{x}$; and a text encoder $\bm{g}_{\bm{\varphi}}:\mathcal{T}\mapsto\mathbb{R}^{N\times 1}$, parameterized by $\bm{\varphi}$, which computes all textual embeddings $\{\bm{g}_{\bm{\varphi}}(\bm{t}_{i})\}_{i=1}^{C}$. Here, $\bm{t}_{i}$ denotes the text template for the $i$-th label, and $\mathcal{T}$ encompasses all candidate text templates. We estimate the raw score in Eq. (1) by
$s_{i}(\bm{x})=\frac{\langle\bm{f}_{\bm{\phi}}(\bm{x}),\bm{g}_{\bm{\varphi}}(\bm{t}_{i})\rangle}{\tau},$ (4)
[…]
IV-C Bi-level Optimization
Our goal is to optimize the joint embedding function, parameterized by $\bm{\phi}$ and $\bm{\varphi}$, which outputs the raw scores for all labels, based on which their marginal probabilities $\hat{p}(y_{i}|\bm{x};\bm{\phi},\bm{\varphi})$ can be computed. Given a training minibatch $\mathcal{B}_{\mathrm{tr}}=\{\bm{x}^{(k)},\bm{y}^{(k)},\mathcal{Z}^{(k)}\}_{k=1}^{K}$, where $\bm{y}^{(k)}=\{0,1\}^{C}$ is the complete ground-truth vector and $\mathcal{Z}^{(k)}\subseteq\{1,\ldots,C\}$ is the index set of observed labels, we minimize an overall loss function that is a linear weighted summation of the individual losses for each observed label:
$\ell(\mathcal{B}_{\mathrm{tr}};\bm{\phi},\bm{\varphi})=$ $\frac{1}{K}\sum_{k=1}^{K}\sum_{i\in\mathcal{Z}^{(k)}}\lambda_{i}\ell\left(\bm{x}^{(k)},y^{(k)}_{i};\bm{\phi},\bm{\varphi}\right),$ (5)
where $\bm{\lambda}=[\lambda_{1},\ldots,\lambda_{C}]^{\intercal}$ is the loss weighting vector to be optimized together with the model parameters $\bm{\phi}$ and $\bm{\varphi}$.
While the de facto cross-entropy function can be used to implement $\ell(\cdot)$ in Eq. (5) for parameter estimation, it is unbounded from above [44]. This can result in excessive penalties for challenging training examples, hindering model capacity allocation and loss weighting adjustment in multitask learning. To address this issue, we resort to the fidelity loss [44], derived from quantum physics, to measure the statistical distance between two states of a quantum:
$\ell(\bm{x},y_{i};\bm{\phi},\bm{\varphi})=$ $1-\sqrt{p(y_{i}|\bm{x})\hat{p}(y_{i}|\bm{x})}$
$-\sqrt{(1-p(y_{i}|\bm{x}))(1-\hat{p}(y_{i}|\bm{x}))},$ (6)
where $p(y_{i}|\bm{x})$ denotes the $i$-th ground-truth label.
To automatically adjust the loss weighting vector $\bm{\lambda}$ and prioritize the primary DeepFake detection task, we employ bi-level optimization [43]. At the outer level, we minimize the primary task concerning $\bm{\lambda}$ on the validation minibatch $\mathcal{B}_{\mathrm{val}}$ to measure and promote generalization. At the inner level, we minimize the overall loss concerning the model parameters $\bm{\phi}$ and $\bm{\varphi}$ on the training minibatch $\mathcal{B}_{\mathrm{tr}}$. This gives rise to the following bi-level optimization problem:
$\begin{split}\min_{\bm{\lambda}}\frac{1}{|\mathcal{B}_{\mathrm{val}}|}&\sum_{\bm{x}\in\mathcal{B}_{\mathrm{val}}}\ell\left(\bm{x},y_{1};{\bm{\phi}}^{\star},{\bm{\varphi}}^{\star}\right)\\ \text{s.t.}\quad{\bm{\phi}}^{\star},{\bm{\varphi}}^{\star}&=\operatorname*{arg\,min}_{\bm{\phi},\bm{\varphi}}\ell(\mathcal{B}_{\mathrm{tr}};\bm{\phi},\bm{\varphi}),\end{split}$ (7)
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. As suggested in [41], we share a common loss weighting for tasks within the same level. Specifically, we learn to adjust three $\lambda_{i}$’s: one for the primary task, one for tasks at the global image attribute level, and one for the local image region level. We sample training and validation data as different minibatches in the same training dataset. We utilize the finite difference method as an efficient approximation to solve Problem (7).
V Experiments
In this section, we first describe the experimental setups and then thoroughly compare the proposed SJEDD with contemporary DeepFake detectors. In addition, we perform a series of ablation experiments to validate the efficacy of SJEDD from both the model and the data perspectives.
V-A Experimental Setups
V-A1 Datasets
We adopt the widely used FF+ [18] and newly established FFSC [41] for training. FF++ contains $1,000$ real videos, manipulated by four DeepFake methods: Deepfakes (DF) [1], Face2Face (F2F) [84], FaceSwap (FS) [83], and NeuralTexures (NT) [4]. These videos come with three compression levels, and our experiments are conducted at the slight compression level, as recommended by [24, 66, 28]. We expand FF++ using the proposed dataset expansion technique described in Sec. III. In contrast, FFSC [41] organizes twelve face manipulations into a semantic hierarchical graph with five global face attribute nodes and six local face region nodes and requires no dataset expansion.
We test DeepFake detectors on four independent DeepFake datasets: CDF [46], FaceShifter (FSh) [45], DeeperForensics-1.0 (DF-1.0) [47], and DeepFake Detection Challenge (DFDC) [48]. Notably, CDF and DFDC are particularly challenging. CDF is difficult due to its high-quality, realistic face forgeries. DFDC poses challenges with fake videos created from real-world footage, often featuring faces in low lighting, with extreme poses, and under occlusions.
V-A2 Implementation Details
SJEDD relies on the CLIP model [42] for joint embedding, where the image encoder is ViT-B/32 [86] and the text encoder is GPT-2 [87]. SJEDD is fine-tuned by minimizing the Problem (7) using AdamW [88] with a decoupled weight decay of $31e-$ 3 and a minibatch size of $32$. The initial learning rate is set to $76e-$ 7, following a cosine annealing schedule [89]. The loss weightings are all initialized to one and optimized using Adam [90] with a learning rate of $31e-$ 3. Throughout training, the temperature parameter $\tau$ in Eq. (4) is fixed at $10$. Following the practices in [25, 19, 27], we randomly introduce one of five real perturbations to each input image with a probability of $0.3$: patch substitution, additive white Gaussian noise, Gaussian blurring, pixelation, and JPEG compression. These perturbations are constrained to be mild to preserve face semantics (i.e., label-preserving). The training process runs for a maximum of $20$ epochs, taking approximately $21$ hours on a single NVIDIA RTX 3090 GPU.
TABLE II: AUC results of the proposed SJEDD against eighteen DeepFake detectors in the cross-dataset setting. All models are trained using the training set of FF++ (and its augmented versions). We retrain CNND, Face X-ray, MADD, FRDM, SLADD, DCL, and SO-ViT-B to match the performance reported in their original papers. For PCL, SFDG, VLFFD, and F2Trans-B, we directly use the results from their corresponding papers due to the lack of publicly available implementations. The last column shows the mean AUC numbers, including/excluding the FF++ test set. The best results are highlighted in bold
Method Venue Backbone FF++ [18] CDF [46] FSh [45] DF-1.0 [47] DFDC [48] FFSC [41] Mean AUC
CNND [19] CVPR2020 ResNet-50 98.13 75.60 65.70 74.40 72.10 61.58 74.59 / 69.88
Face X-ray [21] CVPR2020 HRNet 98.37 80.43 92.80 86.80 65.50 76.57 83.41 / 80.42
PCL [22] ICCV2021 ResNet-34 99.11 81.80 99.40 67.50
MADD [27] CVPR2021 EfficientNet-b4 98.97 77.44 97.17 66.58 67.94 80.91 81.50 / 78.01
Lip-Forensics [28] CVPR2021 ResNet-18 99.70 82.40 97.10 97.60 73.50
FRDM [29] CVPR2021 Xception 98.36 79.40 97.48 73.80 79.70 75.61 84.06 / 81.20
RECCE [23] CVPR2022 Xception 99.32 68.71 70.58 74.10 69.06 77.86 76.61 / 72.06
SBI [30] CVPR2022 EfficientNet-b4 99.64 93.18 97.40 77.70 72.42 78.92 86.54 / 83.92
ICT [31] CVPR2022 ViT-B 90.22 85.71 95.97 93.57 76.74 77.82 86.67 / 85.96
SLADD [24] CVPR2022 Xception 98.40 79.70 93.73 92.80 81.80 76.94 87.23 / 84.99
DCL [32] AAAI2022 EfficientNet-b4 99.30 82.30 97.29 97.48 76.71 74.06 87.86 / 85.57
CADDM [25] CVPR2023 ResNet-34 99.70 91.15 97.32 81.92 71.49 81.22 87.13 / 84.62
SFDG [34] CVPR2023 ResNet-50 95.98 75.83 92.10 73.64
VLFFD [33] ArXiv2023 ViT-L 99.23 84.80
F2Trans-B [35] TIFS2023 99.74 83.05 71.81
MRL [36] TIFS2023 MC3-18 98.27 83.58 92.38 88.81 71.53 76.67 85.21 / 82.59
CFM [37] TIFS2024 EfficientNet-b4 99.62 89.65 93.34 96.38 80.22 78.06 89.55 / 87.53
SO-ViT-B [41] ArXiv2024 ViT-B 98.59 87.53 95.01 91.92 79.58 83.32 89.33 / 87.47
SJEDD (Ours) ViT-B 99.86 92.22 97.80 93.21 84.14 85.70 92.16 / 90.61
V-A3 Competing Methods
We compare SJEDD against eighteen state-of-the-art DeepFake detectors: CNND [19], Face X-ray [21], PCL [22], MADD [27], Lip-Forensics [28], FRDM [29], RECCE [23], SBI [30], ICT [31], SLADD [24], DCL [32], CADDM [25], SFDG [34], VLFFD [33], F2Trans-B [35], MRL [36], CFM [37], and SO-ViT-B [41]. CNND serves as a typical baseline model. MADD employs attention mechanisms for feature selection and aggregation. FRDM and F2Trans-B expose face forgeries through high-frequency analysis. DCL and CFM leverage contrastive learning, while SFDG and MRL utilize graph learning to improve DeepFake detection. Face X-ray, PCL, SLADD, and CADDM incorporate manipulation localization as a manipulation-oriented auxiliary task to facilitate DeepFake detection. SLADD and VLFFD also include $C$-way manipulation classification. RECCE suggests reconstructing face images additionally to learn generative face features. Various empirical losses such as region overlap loss [27], additive margin softmax loss [29], InfoNCE loss [32], and ArcFace loss [31] have also been applied to enhance discriminative feature learning. Furthermore, Face X-ray, PCL, SBI, ICT, SLADD, and VLFFD exploit different data augmentation strategies to enhance detection performance, including face swapping between real faces [21, 22, 30, 31] and manipulation of the local face region [24, 33]. Lip-Forensics and ICT rely on high-level lipreading and face identity features, respectively. Like the proposed SJEDD, SO-ViT-B employs a semantics-oriented multitask learning paradigm but differs in both data and model aspects.
V-A4 Evaluation Metrics
In line with [28, 29, 30, 31, 24, 25, 23, 32, 27, 26], we evaluate the performance of a DeepFake detector using the area under the receiver operating characteristic curve (AUC (%)).
V-B Comparison with Contemporary Detectors
V-B1 Training on FF++
We train SJEDD on the expanded FF++ dataset [18] and evaluate it on the test set of the original FF++ as well as five other DeepFake datasets [46, 45, 47, 48, 41].
Cross-dataset evaluation. Table II shows the AUC results, revealing several interesting observations. First, many detectors struggle with generalization in the cross-dataset setting. In contrast, SJEDD performs consistently well on all five DeepFake datasets, particularly excelling in challenging datasets such as CDF, DFDC, and FFSC, and achieves the highest mean AUC results. Second, all methods show saturated performance on FF++. Third, while SBI reports a high AUC of $93.18\%$ on CDF, it performs poorly on DF-1.0 and DFDC. Similarly, PCL achieves an exceedingly high AUC of $99.40\%$ on DF-1.0 but generalizes poorly on DFDC. This performance discrepancy may be due to heavily relying on manipulation-specific features, such as statistical inconsistencies in color appearances and landmark locations. Fourth, all competing methods score relatively low on DFDC, likely resulting from the significant domain shift caused by different imaging conditions and manipulation techniques. SJEDD achieves an impressive AUC of $84.14\%$ in DFDC, demonstrating the superior transferability of features learned through our semantics-oriented joint embedding paradigm. Last, adhering to the same semantics-oriented design philosophy, SJEDD outperforms its predecessor, SO-ViT-B, confirming the changes in model design (i.e., from discriminative learning to joint embedding) and loss function (i.e., from cross-entropy loss to fidelity loss in bi-level optimization).
TABLE III: AUC results in the cross-manipulation test setting. DF, F2F, FS, and NT represent Deepfakes [1], Face2Face [84], FaceSwap [83], and NeuralTexures [4], respectively. Gray numbers represent within-manipulation results
Training Method DF F2F FS NT Mean AUC
DF MADD [27] 99.51 66.41 67.33 66.01 66.58
RECCE [23] 99.65 70.66 74.29 67.34 70.76
DCL [32] 99.98 77.13 61.01 75.01 71.05
VLFFD [33] 99.97 87.46 74.40 76.79 79.55
CFM [37] 99.93 77.56 54.94 75.04 69.18
SO-ViT-B [41] 99.39 88.25 89.31 82.67 86.74
SJEDD (Ours) 99.51 83.89 95.02 83.28 87.40
F2F MADD [27] 73.04 97.96 65.10 71.88 70.01
RECCE [23] 75.99 98.06 64.53 72.32 70.95
DCL [32] 91.91 99.21 59.58 66.67 72.72
VLFFD [33] 94.90 99.30 65.19 66.69 75.59
CFM [37] 81.85 99.23 60.12 70.80 70.92
SO-ViT-B [41] 99.68 99.39 97.70 92.05 96.47
SJEDD (Ours) 99.83 99.46 97.45 92.34 96.54
FS MADD [27] 82.33 61.65 98.82 54.79 66.26
RECCE [23] 82.39 64.44 98.82 56.70 67.84
DCL [32] 74.80 69.75 99.90 52.60 65.72
VLFFD [33] 92.98 79.69 99.57 53.53 75.40
CFM [37] 72.91 71.39 99.85 51.69 65.33
SO-ViT-B [41] 98.88 76.48 96.07 75.38 83.58
SJEDD (Ours) 99.68 81.84 93.29 81.19 87.57
NT MADD [27] 74.56 80.61 60.90 93.34 72.02
RECCE [23] 78.83 80.89 63.70 93.63 74.47
DCL [32] 91.23 52.13 79.31 98.97 74.22
VLFFD [33] 92.74 62.53 85.62 98.99 80.30
CFM [37] 88.31 76.78 52.56 99.24 72.55
SO-ViT-B [41] 99.40 94.86 94.25 95.11 96.16
SJEDD (Ours) 99.88 95.14 97.31 98.51 97.46
Cross-manipulation evaluation. We perform the cross-manipulation evaluation of SJEDD against six representative DeepFake detectors: MADD [27], RECCE [23], DCL [32], VLFFD [33], CFM [37], and SO-ViT-B [41]. Specifically, we train on one manipulation type from FF++ [18] (including additional augmented data) and test on the remaining manipulations from FF++. As shown in Table III, SJEDD surpasses other methods by clear margins, which we believe arises from two design choices. First, our semantics-oriented formulation helps SJEDD learn generalizable features across different manipulations that alter the same face attributes (e.g., NT $\rightarrow$ F2F, both modifying the global expression attribute and the local mouth and lip regions). Second, SJEDD models the joint probability distribution of face attributes (see Eq. (1)), thus promoting generalization across different face attributes (e.g., DF: identity $\rightarrow$ NT: expression). Once again, SJEDD outperforms its predecessor SO-ViT-B, highlighting the effectiveness of the joint embedding formulation and the fidelity loss in bi-level optimization.
Cross-attribute evaluation. As suggested in FFSC [41], we adopt two semantics-oriented testing protocols: Protocol-1, which focuses on generalization to novel manipulations for the same face attribute, and Protocol-2, which examines generalization to novel face attributes. Six DeepFake detectors are chosen for comparison, including Lip-Forensics [28], SLADD [24], ICT [31], DCL [32], CFM [37], and SO-ViT-B [41]. According to Table IV, SLADD, DCL, and CFM perform poorly in the Protocol-1 test. This suggests that their learned features are manipulation-specific, thus falling short of detecting the same face attribute forgeries when altered by different methods. In stark contrast, semantics-oriented methods such as SO-ViT-B and the proposed SJEDD yield favorable results in the Protocol-1 test, underscoring the advantages of learning and integrating high-level semantic features at the global face attribute level with low-level signal features at the local face region level. When switching to the Protocol-2 test, SJEDD continues to outperform manipulation-oriented methods, particularly in generalizing to the unseen gender attribute. This success may be attributed to the learned transferable semantic and signal features from various local face regions and across different global face attributes, thereby facilitating generalization to novel face attributes.
TABLE IV: AUC results in the cross-attribute test setting. All models are trained on the training set of FF++ (and its augmented versions). “–” indicates that no such evaluation can be properly conducted
Method Intra-dataset Protocol-1 Protocol-2 Mean AUC
Expression Identity Age Gender Pose
Lip-Forensics [28] 99.70 76.76 60.75 76.76
SLADD [24] 98.40 82.39 65.55 85.31 81.21 70.00 76.89
ICT [31] 90.22 84.20 72.18 52.83 76.12
DCL [32] 99.30 65.72 85.25 84.11 75.41 71.67 76.43
CFM [37] 99.62 73.37 89.97 86.76 77.68 74.59 80.47
SO-ViT-B [41] 98.59 76.89 92.83 86.39 81.25 88.42 85.16
SJEDD (Ours) 99.86 80.16 92.23 86.45 86.37 89.03 86.87
TABLE V: AUC results of SJEDD against six contemporary DeepFake detectors in the cross-dataset test setting. All models are trained on FFSC
Method FF++ CDF DF-1.0 DFDC Mean AUC
CNND [19] 80.57 83.93 85.98 72.99 80.87
MADD [27] 87.63 87.46 83.96 77.02 84.02
FRDM [29] 89.07 80.39 88.46 73.58 82.87
RECCE [23] 95.56 78.51 69.73 67.20 77.77
CADDM [25] 83.29 81.10 87.04 65.23 79.17
SO-ViT-B [41] 86.33 88.43 92.46 79.57 86.77
SJEDD (Ours) 93.81 91.28 97.60 81.32 91.00
V-B2 Training on FFSC
We next train the proposed SJEDD on FFSC [41], which encompasses five global face attributes—age, expression, gender, identity, and pose—instantiated by twelve face manipulations. The connections between global face attribute nodes and local face region nodes, as well as the manipulation degree parameters, are determined through formal psychophysical experiments [41]. The text templates to embed the label hierarchy are identical to those described in Sec. IV-B. We conduct the cross-dataset evaluation of SJEDD against six DeepFake detectors: CNND [19], MADD [27], FRDM [29], RECCE [23], CADDM [25], and SO-ViT-B [41] on FF++ [18], CDF [46], DF-1.0 [47], and DFDC [48]. All competing methods are trained on FFSC.
As shown in Table V, SJEDD performs significantly better than the competing detectors on the four unseen DeepFake datasets. This shows the effectiveness and generality of the proposed semantics-oriented joint embedding in enhancing DeepFake detection across various training datasets. Additionally, all detectors demonstrate strong performance on FF++, with RECCE achieving an impressive AUC of $95.56\%$. This can be attributed to the relatively small domain gap between FF++ and FFSC, as a portion of images in these datasets share the same video source (i.e., YouTube). Consequently, RECCE can easily “shortcut” by learning signal-level features through face reconstruction to improve DeepFake detection on FF++. In contrast, the detection accuracy on DFDC is comparatively lower, consistent with the observations in Table II.
V-C Ablation Studies
In this subsection, we conduct a series of ablation experiments to verify SJEDD’s design choices from both the model and data perspectives.
V-C1 Ablation from the Model Perspective
We first analyze two key components of SJEDD: the joint embedding in Eq. (4) and the fidelity loss in Eq. (IV-C). To assess their effectiveness, we construct two SJEDD degenerates: one implemented solely by the image encoder to produce the raw scores (i.e., $\bm{s}(\bm{x})=\bm{f}_{\bm{\phi}}(\bm{x})$) and trained with the cross-entropy loss, and the other implemented by joint embedding and trained with the cross-entropy loss. Table VI shows the cross-dataset evaluation results, in which all model variants are trained on FFSC.
TABLE VI: AUC results of two degenerates of SJEDD. The term “w/o joint embedding” refers to directly predicting the label hierarchy using only the image encoder. “w/o fidelity loss” means substituting the fidelity loss with the cross-entropy loss during training. All models are trained on FFSC
Joint Embedding Fidelity Loss FF++ CDF DF-1.0 DFDC Mean AUC
86.63 88.43 92.46 79.57 86.77
91.77 90.23 96.22 81.35 89.89
93.81 91.28 97.60 81.32 91.00
It is clear from the table that joint embedding achieves noticeably better performance than directly predicting the label hierarchy from the input face image. This highlights the effectiveness of joint embedding in multitask learning, where the model can automatically adjust its capacity for each task through end-to-end optimization. Additionally, joint embedding facilitates task relationship modeling as distances between text embeddings in the learned feature space. By further incorporating the loss of fidelity, SJEDD achieves the best results. As discussed previously, this could be due to the favorable mathematical properties of the fidelity loss compared to the cross-entropy loss [44]. Fig. 4 illustrates the training dynamics of these two loss functions, where we find that the cross-entropy loss exhibits pronounced fluctuations, whereas the fidelity loss curve is more stable. This provides additional evidence that the fidelity loss may guide the optimization towards better local optima.
Figure 4: Training dynamics of the fidelity loss versus the cross-entropy loss.
(a) FFSC: Age
(b) FFSC: Expression (Smile)
(c) FFSC: Expression (Surprise)
(d) FFSC: Gender
(e) FFSC: Identity
(f) FFSC: Pose
(g) FF++: DF
(h) FF++: F2F
(i) FF++: FS
(j) FF++: NT
(k) CDF
(l) DFDC
Figure 5: Illustration of the proposed SJEDD trained on FFSC in making predictions at the face attribute and region levels. For FFSC (as an intra-dataset test), fake attributes/regions are highlighted in bold. For FF++, CDF, and DFDC (as cross-dataset tests), only the inference results are provided for reference. Zoom in for improved visibility.
[…]
$s_{i}(\bm{x})=\frac{\langle\bm{f}_{\bm{\phi}}(\bm{x}),\bm{g}_{\bm{\varphi}}(\bm{t}_{i})\rangle}{\tau\|\bm{f}_{\bm{\phi}}(\bm{x})\|_{2}\|\bm{g}_{\bm{\varphi}}(\bm{t}_{i})\|_{2}},$ (8)
where $\|\cdot\|_{2}$ denotes the $\ell_{2}$-norm and $\tau$ is the same temperature parameter as in Eq. (4), and 5) optimizing SJEDD with a fixed set of loss weightings $\{1.0,0.1,0.1\}$ in Eq. (5) to prioritize the primary task manually.
TABLE VII: AUC results of different variants of SJEDD, all trained on the expanded FF++
Model Variant CDF DF-1.0 DFDC FFSC Mean AUC
w/ Frozen $\bm{f}_{\bm{\phi}}$ + $\bm{g}_{\bm{\varphi}}$ 50.00 50.00 50.00 50.00 50.00
w/ Frozen $\bm{g}_{\bm{\varphi}}$ 92.38 91.98 81.54 83.08 87.25
w/ Simplifying Encoding 90.49 92.15 83.83 85.33 87.95
w/ Normalized Similarity 89.00 93.15 84.01 83.96 87.53
w/o Bi-level Optimization 88.90 92.04 79.94 83.75 86.16
SJEDD (Ours) 92.22 93.21 84.14 85.70 88.82
Table VII presents the cross-dataset results with several useful findings. First, freezing both the image encoder and the text encoder (i.e., using the pre-trained CLIP for zero-shot DeepFake detection) proves ineffective in this task. Fine-tuning the image encoder significantly boosts the detection performance, which validates the importance of vision-language alignment. Joint fine-tuning of the image and text encoders offers additional performance gains.
Additionally, SJEDD is fairly robust to changes in the format of the text templates, which experiences only slight performance drops when using simplified keyword-based templates. Using normalized similarity for raw score calculation negatively impacts the detection performance, indicating that some discriminative information may be captured in the magnitude of the visual embedding. Finally, dynamically adjusting the loss weightings and prioritizing the primary task through bi-level optimization proves advantageous, which also removes the need for cumbersome hyperparameter tuning.
V-C2 Ablation from the Data Perspective
To evaluate the effectiveness of the proposed semantics-oriented dataset expansion technique, we first isolate the roles of its two steps: data augmentation and data relabeling. We then compare our method with the widely used Face X-ray data augmentation [21] and the human-involved data annotation in FFSC [41]. Tables VIII and IX present the results, where models are trained on FF+ [18] (including any dataset expansion if applicable) by default unless stated otherwise.
TABLE VIII: AUC results of SJEDD trained using different variants of the proposed semantics-oriented dataset expansion technique applied to FF++. “w/o relabeling” corresponds to the binary classification paradigm
Augmentation Relabeling CDF DF-1.0 DFDC FFSC Mean AUC
81.45 86.78 76.00 78.04 80.57
86.39 91.07 81.39 82.52 85.34
84.58 87.37 79.19 78.56 82.43
92.22 93.21 84.14 85.70 88.82
As shown in Table VIII, data relabeling alone marginally improves SJEDD’s generalization to unseen datasets. This happens because, in FF++, the relabeling process simply merges DF [1] and FS [83] into one category, and F2F [84] and NT [4] into another, which does not entirely fulfill the potential of semantics-oriented categorization of face manipulations. As a result, SJEDD reduces to a manipulation-oriented detector with a risk of overfitting. Training with data augmentation consistently improves detection generalizability, aligning with the findings in [21, 24, 30]. The best performance is achieved by the complete SJEDD, in which both data augmentation and relabeling are activated. The improvements arise because our proposed semantics-oriented dataset expansion method enables a single manipulation to alter multiple attributes and meanwhile group multiple manipulations under the same attribute category. This encourages SJEDD to learn transferable features across different manipulations that affect similar face attributes, rather than relying solely on manipulation-specific cues.
(a) FF++: DF
(b) FF++: F2F
(c) FF++: FS
(d) FF++: NT
(e) CDF
(f) FSh
(g) DF-1.0
(h) DFDC
Figure 6: Visualization of SJEDD activation maps. In each subfigure, the forged face image is displayed on the left, with the corresponding attention heatmap on the right. Best viewed in color.
TABLE IX: AUC results of SJEDD variants trained using different data augmentation methods
Training FF++ CDF DF-1.0 DFDC FFSC
Face X-ray-augmented FF++ 89.58 93.28 81.14 82.15
Human-annotated FFSC 93.81 91.28 97.16 81.32
Semantics-oriented FF++ (Ours) 92.22 93.21 84.14 85.70
As shown in Table IX, the Face X-ray-augmented SJEDD performs worse than the proposed SJEDD. Furthermore, when tested on CDF [46], DF-1.0 [47], and DFDC [48], the proposed SJEDD achieves results comparable to the FFSC-trained SJEDD, which is exposed to a larger set of real-world sophisticated face manipulations and more comprehensive global face attributes. This underscores the effectiveness of our method in automatically expanding existing DeepFake datasets and promoting their generality. Additionally, models trained on FF++ do not generalize well to FFSC compared to models trained on FFSC and tested on FF++. This suggests that FFSC presents a more challenging test set.
V-D Interpretability
The proposed SJEDD via vision-language correspondence enhances model interpretability by providing human-understandable explanations. Specifically, we leverage SJEDD trained on FFSC [41] to illustrate this interpretability, as shown in Fig. 5. SJEDD accurately highlights the modified face attributes and regions with high probabilities. For example, in Fig. 5 (b), the detector identifies the face as fake and meanwhile indicates a likely expression alteration, pinpointing the mouth and lip as manipulated regions. This allows us to combine the predefined text templates and form a textual explanation such as “The face image is fake because the global face attribute of expression is altered, through manipulation of the local lip and mouth regions.” Regarding the cross-dataset results on FF++, CDF, and DFDC, SJEDD can still successfully interpret the binary classification results by predicting the authenticity of individual face attributes and regions. For example, on FF++, our detector accurately identifies DF and FS as instances of identity manipulation and F2F and NT as examples of expression forgery.
We also present SJEDD activation maps by Grad-CAM [91] in Fig. 6, where warmer regions indicate more significant importance. As can be seen, SJEDD consistently focuses on the face regions to identify forgeries, aligning with our primary design goal of semantics-oriented DeepFake detection.
VI Conclusion and Discussion
We have described a semantics-oriented joint embedding DeepFake detection method called SJEDD, along with a dataset expansion technique. The core of SJEDD lies in its use of joint embedding, which reformulates model parameter splitting in multitask learning to a model capacity allocation problem, allowing for end-to-end bi-level optimization. We replaced the cross-entropy loss with the fidelity loss in bi-level optimization, which additionally boosts the performance. SJEDD also renders some degree of model interpretation by providing human-understandable explanations. We hope that the development of SJEDD will inspire researchers in related fields to rethink DeepFake detection from a semantic perspective.
The present study addresses DeepFake detection purely from a technical standpoint. Future directions in this field are multifaceted, involving advancements in both technology and policy. From a technological perspective, further exploration of a multimodal approach, which combines audio, visual, and textual analysis, is recommended to enhance detection accuracy. Detectors should also be able to operate in real-time on edge devices to support live analysis of streaming content in large-scale media platforms. Another important avenue for advancement is to develop detectors capable of continual learning and adaptation to emerging DeepFake techniques, potentially involving human oversight to maintain detection accuracy and efficacy. Proactive measures such as using blockchain technology and digital watermarking to authenticate original content should also be considered. From a policy perspective, it is essential to establish international standards and frameworks for DeepFake detection. This would involve collaboration between governments, tech companies, and research institutions. Laws and regulations should also be enacted to address the creation and distribution of DeepFake content, including penalties for misuse and support for victims.